# Examining Long-Context Large Language Models for Environmental Review Document Comprehension

### v2

@@
Abstract
As LLMs become increasingly ubiquitous, researchers have tried various techniques to augment the knowledge provided to these models. Long context and retrieval-augmented generation (RAG) are two such methods that have recently gained popularity. In this work, we examine the benefits of both of these techniques by utilizing question answering (QA) task in a niche domain. While the effectiveness of LLM-based QA systems has already been established at an acceptable level in popular domains such as trivia and literature, it has not often been established in niche domains that traditionally require specialized expertise. We construct the NEPAQuAD1.0 benchmark to evaluate the performance of five long-context LLMs – Claude Sonnet, Gemini, GPT-4, Llama 3.1, and Mistral – when answering questions originating from Environmental Impact Statements prepared by U.S. federal government agencies in accordance with the National Environmental Environmental Act (NEPA). We specifically measure the ability of LLMs to understand the nuances of legal, technical, and compliance-related information present in NEPA documents in different contextual scenarios. We test the LLMs’ internal prior NEPA knowledge by providing questions without any context, as well as assess how LLMs synthesize the contextual information present in long NEPA documents to facilitate the question/answering task. We compare the performance of the models in handling different types of questions (e.g., problem-solving, divergent, etc.). Our results suggest that RAG powered models significantly outperform those provided with only the PDF context in terms of answer accuracy, regardless of the choice of the LLM. Our further analysis reveals that many models perform better answering closed type questions (Yes/No) than divergent and problem-solving questions.
1 Introduction
As Large Language Model (LLM) s become increasingly commonplace, researchers have discovered that these models are useful for many tasks beyond text generation. Specifically, LLM s have shown potential utility in niche domains (like science) that would traditionally require specialized expertise, both in a pure text setting Horawalavithana et al. (2022); Munikoti et al. (2024) and by incorporating data of various modalities Dollar et al. (2022); Horawalavithana et al. (2023). Recent work has been done to evaluate these models Acharya et al. (2023); Munikoti et al. (2023); Cai et al. (2024) and to assess their uncertainty Wagle et al. (2024). Despite extensive research, constructing LLM s for answering domain-specific questions has proven challenging Kasneci et al. (2023).
One such challenge for LLM-based question-answering systems occurs when systems are tasked with surfacing answers to questions from the content of long documents in specialized domains. Existing LLM s allow users to include a paragraph as context along with the content of the question; however, LLM s generally limit the size of that paragraph to a specific number of tokens. This restriction forces users to truncate or manually summarize the content of lengthy documents into short passages. Another approach users can take includes submitting only the question and relying on the models to find the correct document from a corpus and relevant content needed to answer the question. This strategy often works well for answering questions from well-known domains (e.g., sports or education); however, it is not as successful for less pervasive topics Munikoti et al. (2023). Because LLM s are data-driven, they are not as apt to provide accurate answers for questions about less popular, more specialized domains such as Law Kapoor et al. (2024)) and Energy Buster et al. (2024).
In this work, we focus on assessing the long-context LLM s in the environmental reviews conducted under the National Environmental Protection Act (NEPA)  [^1] https://www.epa.gov/nepa. NEPA is a U.S. environmental law designed to protect the environment. An environmental impact statement (EIS) is required by Section 102(2)(C) of NEPA for any proposed major federal action significantly affecting the quality of the environment. An EIS is a detailed document prepared that describes a proposed action, alternatives to the proposed action, and potential effects of the proposed action and alternatives on the environment. An EIS contains information about environmental permitting and policy decisions and considers a range of reasonable alternatives, analyzes the potential impacts resulting from the proposed action and alternatives, and demonstrates compliance with other applicable environmental laws and executive orders.
Along with the fact that EIS documents are usually lengthy (often several hundred pages) and are created by NEPA experts, another factor that can hinder the application of LLM s in this domain is that the development of an EIS document requires NEPA experts with various subject matter expertise to engage in preparation over multiple years, often citing older articles from as far back as the 1990s. For example, the Executive Order (EO) 12898, issued in 1994, is cited on page 60 of the EIS documents prepared for the First Responder Network Authority project [^2] https://www.energy.gov/nepa/eis-0530-nationwide-public-safety-broadband-network-programmatic-environmental-impact. This could present significant challenges for current LLM s in helping NEPA users automatically retrieve answers from LLM-based question-answering systems. To our knowledge, there is no ground-truth benchmark built specifically for this domain to evaluate the quality of LLM s’ output for QA task when the questions pertain to EIS documents.
In this work, we leverage both long-context LLM s and Retrieval Augmented Generation (RAG) techniques to evaluate the extent to which LLM s can comprehend NEPA knowledge for question-answering (Figure 1). We select frontier LLM s for our experiments: Claude Sonnet Anthropic Team and Collaborators (2024), Gemini Gemini Team and Collaborators (2024), GPT-4 OpenAI (2024), Llama 3.1Dubey et al. (2024) and MistralJiang et al. (2023a). We also conduct rigorous experiments evaluating LLM s with various types of contexts for NEPA documents. To evaluate our approach, we establish a benchmark using ground truth triples of questions, answers, and corresponding contexts, generated through a semi-supervised method employing GPT-4. Overall, we make the following contributions:
  1. 1.
    Created the first-ever preliminary benchmark (NEPAQuAD1.0) to automatically evaluate the performance of LLM s in a question-answering task for EIS documents
  2. 2.
    Evaluated the capability of LLM s in question-answering tasks over long documents
  3. 3.
    Conducted rigorous comparisons of LLM s using zero-shot prompting versus context-driven prompting (i.e., passage, PDF, and RAG) to assess model performance.
The structure of this paper is outlined as follows: In Section 2, we describe the benchmark creation to assess the quality of our model in comparison to models derived from different contexts. The Section 3 section lays out our approach and the various contexts used for evaluation implementation, followed by a detailed analysis of our performance in Sections 4. Section 5 then discusses other works in literature that deal with long context and RAG for long documents. We finish with the conclusion and limitations of our work in Sections 6 and 7.
Figure 1: Illustration of varied EIS contexts used in the comparison.
2 NEPAQuAD Benchmark
In this section, we describe the construction of a ground-truth benchmark to evaluate the quality of automated responses generated from LLM s. Due to the high costs associated with manually creating human-written questions and answers, and the inability to use ground-truth benchmarks from other domains, we adopt a semi-automatic approach to generate the NEPAQuAD1.0 (National Environmental Policy Act Question and Answering Dataset) benchmark. The general idea of our evaluation benchmark generation process is to extract meaningful passages from a set of EIS documents, then use GPT-4 to generate questions based on these passages. To ensure the quality of the generated benchmark, two authors of this study, who are subject matter experts in NEPA, measure the quality of the ground-truth answers by comparing the provided proofs against the original context from which the questions were derived. Our generated ground-truth benchmark is a set of triples containing a question, an answer, and the proof (i.e., the text directly related to the answer, derived from the context from which the question originated). The process of benchmark generation is illustrated in Figure 2.
@@ 2 NEPAQuAD Benchmark
Figure 2: Steps of Ground-truth benchmark generation for evaluating LLM s over varied contexts for question-answering over EIS documents
To evaluate performance of LLM s for the EIS question-answering task, we first select high-quality documents from the EIS document database and extract paragraphs as context to be used in the evaluation. Then, we identify the types of questions that we want to use to evaluate the models. Next, we use GPT-4 OpenAI (2024) to generate question-answer pairs based on the selected contexts by using carefully designed prompts. Finally, we use these generated questions to evaluate different LLM s with various contexts, with the generated answers acting as the ground-truth. We describe the process in detail below.
2.1 Gold Passage Selection
@@
Document Selection NEPA experts select nine EIS documents from different government agencies that were most representative of various NEPA actions. These document exhibit great variations in content and focus depending on the authoring government agency, as each agency may interpret and implement the NEPA guidelines distinctively. For instance, the U.S. Forest Service might emphasize forest management and wildfire mitigation, while the Army Corps of Engineers could prioritize water resource development and infrastructure impacts. Table 3 shows the statistics about the selected documents (see Appendix). Each document has around 400 pages on average while the longest document contain more than 600K tokens.
@@
Excerpt Selection For each of nine selected documents, we attempt to select excerpts that have important content of each document. Again, the default approach of randomly extracting excerpts poses the risk of resulting in low-quality excerpts, such as parts of appendix or images’ captions. To avoid this risk, NEPA experts manually select excerpts from the documents. They divided each document into three sections: beginning, middle, and end, and then selected two, six, and two excerpts from each of these sections respectively, for a total of 10 excerpts from each document. We use these excerpts as the ground-truth context, called gold passages, for question benchmark generation.
2.2 Question Type Selection
@@
Once we identified the gold passages, NEPA experts select the type of questions to include in the benchmark. We started with a list of 15 types of questions [^3] https://tinyurl.com/3akej8ct, and eventually narrowed it down to 10 types of questions after extensive discussions. These types are shown in Table 1. In addition to selecting the question types, the NEPA experts also created sample questions for each type. For a more detailed description of the question types, as well as example questions, please see Appendices A and B.
Question Type #Questions
Closed 789 49%
Comparison 64 4%
Convergent 109 7%
Divergent 121 8%
Evaluation 64 4%
Funnel 127 8%
Inference 101 6%
Problem-solving 11 1%
Process 108 7%
Recall 105 7%
Total 1599 100%
Table 1: Statistics on question types used in the NEPAQuAD1.0 benchmark
2.3 Prompt Design
The next step is to design a prompt that can instruct the question generation model to create high quality questions and answers. To ensure that the prompt can instruct generative model efficiently, we took advice from the NEPA experts to create the prompts. We also use the sample questions created by the NEPA experts to augment the original prompts and create an “enhanced" prompt. The template for the prompt and benchmark creation process is displayed in Figure 2. We restricted the output for each prompt in a CSV format with three fields: question, answer, and proof. The "proof" column stored the part of the gold passage that the model picked as evidence for the provided answer to the question.
2.4 Ground-truth Benchmark Generation
Automated Generation We selected GPT-4 as the generative model for this task, as GPT-4 has been used for generating questions and answers for documents in other domains such as agriculture Balaguer et al. (2024). Specifically, we use GPT-4 Azure OpenAI service with default setting to execute the question generation prompts. For each document in our nine selected documents, we have 10 gold passages, results in 90 gold passages in total. For each gold passage, we generated 10 sets of closed questions and two each of the other nine types of questions. Out of these total 2520 generated questions, a significant number of these were not outputted in the correct format by the LLM s. We then filtered the generations that were not in the desired formats, bringing the number down to 1920. Finally, we manually filtered the remaining questions for quality, thus ending with a benchmark of 1599 ground-truth triples of question-answer-proof over the nine selected EIS documents. We make the benchmark publicly available. (PolicyAI, 2024).
Quality Check In order to judge the generated benchmark’s correctness, we randomly select 100 sample questions for validation. The validation was done by the NEPA experts (two co-authors). Each of them independently went through the sample questions, checking both the correctness of question type (i.e. whether the generated question was the same type of question as requested) and the correctness of the answer. For each question, if either of the evaluators marked the question or answer as correct, we labelled them as correct. Overall, our generated benchmark achieved $77\%$ of answer correctness and $82\%$ of question type correctness. While these human checks certainly helped assure the quality, we nevertheless still have some limitations in the dataset creation process, as discussed in Section 7.
3 Experimental Setup
Our experimental framework is designed to comprehensively evaluate the performance of various LLM s in responding to questions related to NEPA documents. We selected a diverse set of models, including both closed and open-source options, to provide a broad perspective on current LLM capabilities in this domain. The models under evaluation are: GPT-4 (OpenAI) OpenAI (2024), Claude-3 Sonnet (Anthropic) Anthropic Team and Collaborators (2024), Gemini (Google) Gemini Team and Collaborators (2024), 8B LLAMA3.1 (Meta)Dubey et al. (2024) and 7B Mistral (Mistral AI) Jiang et al. (2023a). To assess these models’ performance across different scenarios, we developed a benchmark consisting of questions derived from NEPA documents. Each model will be evaluated under four distinct context conditions, allowing us to analyze how different levels of information availability affect their responses. The comprehensive evaluation framework will allow us to draw meaningful conclusions about the strengths and limitations of different LLM s in handling NEPA-related queries under varying levels of contextual information.
3.1 Context Variation
We designed four distinct context scenarios that simulate various real-world information retrieval situations. These scenarios are designed to test the models’ capabilities under different levels of available context:

### v3

# Benchmarking LLMs for Environmental Review and Permitting

@@
Abstract.
The National Environment Policy Act (NEPA) stands as a foundational piece of environmental legislation in the United States, requiring federal agencies to consider the environmental impacts of their proposed actions. The primary mechanism for achieving this is through the preparation of Environmental Assessments (EAs) and, for significant impacts, comprehensive Environmental Impact Statement (EIS). Large Language Model (LLM) s’ effectiveness in specialized domains like NEPA remains untested for adoption in federal decision making processes. To address this gap, we present NEPA Question and Answering Dataset (NEPAQuAD), the first comprehensive benchmark derived from EIS documents, along with a modular and transparent evaluation pipeline MAPLE  to assess LLM performance on NEPA focused regulatory reasoning tasks. Our benchmark leverages actual EIS documents to create diverse question types, ranging from factual to complex problem-solving ones. We built a modular and transparent evaluation pipeline to test both closed- and open-source models in zero-shot or context-driven QA benchmarks. We evaluate five state-of-the-art LLM s – Claude Sonnet 3.5, Gemini 1.5 Pro, GPT-4, Llama 3.1, and Mistral-7B-Instruct – using our framework to assess both their prior knowledge and their ability to process NEPA-specific information. The experimental results reveal that all the models consistently achieve their highest performance when provided with the gold passage as context. While comparing the other context-driven approaches for each model, Retrieval Augmented Generation (RAG)-based approaches substantially outperform PDF document contexts indicating that neither model is well suited for long-context question-answering tasks. Our analysis suggests that NEPA focused regulatory reasoning tasks pose a significant challenge for LLM s, particularly in terms of understanding the complex semantics and effectively processing the lengthy regulatory documents.
large language models, retrieval augmented generation, benchmarks, evaluation, environmental permitting
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 03–07, 2025; Toronto, CAisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
LLMs are demonstrating increasingly sophisticated cognitive abilities, including idea generation (Ege et al., 2024), problem-solving (Rasal and Hauer, 2024), and the potential to simulate believable human behaviors such as reflection and planning (Park et al., 2023). These capabilities are enabling LLMs to assist significantly in professional workflows and data-driven decision-making across various sectors, notably in Law (Siino et al., 2025) and Medicine (Kim et al., 2024a). For example, LLMs are already assisting with tasks such as drafting documents, reviewing information, and generally speeding up rote, repetitive, or data-intensive tasks within the decision making processes that are prone to human error or consume significant time (Guha et al., 2023; Ma et al., 2024). However, despite this demonstrated potential and application in other professional domains, their widespread adoption in government decision-making processes remains limited, primarily due to concerns surrounding trust and reliability.
In this work, we assess LLM performance in the domain of environmental reviews conducted under the National Environment Policy Act (NEPA  [^1] https://www.epa.gov/nepa). NEPA stands as a foundational piece of environmental legislation in the United States, requiring federal agencies to consider the environmental impacts of their proposed actions. Beyond simply protecting the environment, the NEPA process plays a important role in promoting societal progress and responsible economic growth. By assuring informed decision-making, promoting transparency, and incorporating public input, NEPA helps shape federal decisions that are more sustainable, resilient, and responsive to community needs. This proactive approach can prevent costly environmental impacts, facilitate the development of necessary infrastructure in an environmentally sound manner, and build public trust, ultimately contributing to long-term economic stability and community well-being.
However, the NEPA decision making process is inherently complex, and characterized by different agency-wide regulations, multi-disciplinary technical analysis, and a deeply collaborative process involving diverse stakeholders including applicants, consultants, regulatory agencies, legal teams, and the public [^2] https://www.epa.gov/nepa/national-environmental-policy-act-review-process. To perform such decisions, it requires more than just comprehensive knowledge of federal, state, and local laws; it demands the ability to interpret their application to specific site conditions, analyze complex environmental interactions, predict potential outcomes, and communicate findings clearly and effectively. NEPA decision makers want the LLMs to perform the higher-order regulatory focused reasoning beyond fact retrieval for effective decision support. To support the adoption of LLMs in NEPA decision making workflows [^3] https://www.whitehouse.gov/presidential-actions/2025/04/updating-permitting-technology-for-the-21st-century/, we test the LLM’s abilities to perform regulatory focused reasoning to interpret the intent behind regulations, apply their principles effectively, and support logical deductions or evaluations grounded in NEPA regulatory frameworks. This involves evaluating potential compliance pathways, identifying likely environmental impacts based on regulatory criteria, recommending appropriate mitigation strategies, and predicting plausible agency review outcomes based on project characteristics.
@@ 1. Introduction
Figure 1. Steps of Ground-truth benchmark generation for evaluating LLM s over varied contexts for question-answering over EIS documents
An environmental impact statement (EIS) is required by Section 102(2)(C) of NEPA for any proposed major federal action significantly affecting the quality of the environment. An EIS is a detailed document prepared that describes a proposed action, alternatives to the proposed action, and potential effects of the proposed action and alternatives on the environment. An EIS contains information about environmental permitting and policy decisions and considers a range of reasonable alternatives, analyzes the potential impacts resulting from the proposed action and alternatives, and demonstrates compliance with other applicable environmental laws and executive orders. Along with the fact that EIS documents are usually lengthy (often several hundred pages) and are created by NEPA experts, another factor that can hinder the application of LLM s in this domain is that the development of an EIS document requires NEPA experts with various subject matter expertise to engage in preparation over multiple years, often citing older articles from as far back as the 1990s. For example, the Executive Order (EO) 12898, issued in 1994, is cited on page 60 of the EIS documents prepared for the First Responder Network Authority project [^4] https://www.energy.gov/nepa/eis-0530-nationwide-public-safety-broadband-network-programmatic-environmental-impact. This could present significant challenges for current LLM s in helping NEPA users automatically retrieve answers from LLM-based question-answering systems. To our knowledge, there is no ground-truth benchmark built specifically for this domain to evaluate the quality of LLM s’ output for QA task when the questions pertain to EIS documents.
To this end, we evaluate the extent to which LLM s can comprehend NEPA knowledge for diverse question typologies specifically designed to assess NEPA focused regulatory reasoning over EIS. We select frontier LLM s for our experiments: Claude Sonnet (Anthropic Team and Collaborators, 2024), Gemini-1.5Pro (Gemini Team and Collaborators, 2024), GPT-4 (OpenAI, 2024), Llama 3.1(Dubey et al., 2024) and Mistral-7B-Instruct(Jiang et al., 2023). We built NEPA Question and Answering Dataset (NEPAQuAD) benchmark that includes the triplets of questions, answers, and corresponding contexts, generated through a hybrid approach employing GPT-4 and NEPA experts. To support our evaluation, we developed a Multi-context Assessment Pipeline for Language model Evaluation (MAPLE), a transparent and modular evaluation pipeline that seamlessly handles response generation across different LLM providers. MAPLE also supports various evaluation scenarios – basic question-answering with no context, and advanced features like RAG-based response generation and context-grounded evaluation. Using the MAPLE framework, we conduct rigorous experiments evaluating frontier LLM s with various types of contexts for NEPA documents. Overall, we make the following contributions:
  1. (1)
    Constructed the first-ever benchmark NEPAQuAD v1.0 to evaluate the performance of LLM s for diverse question typologies specifically designed to assess NEPA focused regulatory reasoning.
  2. (2)
    Developed MAPLE as a standardized evaluation pipeline to compare the capabilities of LLM s in various context-driven prompting strategies (i.e., gold passage, entire PDF document, and RAG) to assess model performance.
  3. (3)
    NEPA experts driven performance evaluation in the LLM s ability to support real-world regulatory decisions.
The structure of this paper is outlined as follows. In Section 2, we describe the NEPAQuAD v1.0 benchmark in detail. Section 3 lays out our approach and the various contexts used for evaluation implementation, followed by a detailed analysis of our performance in Section 4. Section 5 then discusses other socio-scientific benchmarks and evaluation frameworks in the literature. We finish with the conclusion and limitations of our work in Sections 6 and 7.
2. The NEPAQuAD v1.0 Benchmark
We present NEPAQuAD v1.0 (National Environmental Policy Act Question and Answering Dataset Version 1.0), the first-ever benchmark to evaluate the performance of LLMs in a question-answering task for EIS documents. NEPAQuAD consists of 1590 questions, split into two different question types: open and closed. The open questions are further split into nine different types, as shown in Table 1. Overall, these questions not only require LLM s to process and comprehend long-context documents, but also reason in the field of environmental permitting. As such, these questions are perfectly suited to serve as the test for any LLM’s capability to reason in this regulatory domain.
2.1. Dataset Creation
Due to the high costs associated with manually creating the entire dataset and the inability to use ground-truth benchmarks from other domains, we adopt a hybrid approach with GPT-4 and NEPA experts to generate the NEPAQuAD v1.0 benchmark. The process we followed for this benchmark generation is illustrated in Figure 1, which we describe in more detail in the following sections:
@@ 2.1. Dataset Creation
Document Curation The first task is to curate a diverse set of high-quality EIS documents from federal databases and select excerpts to establish a representative evaluation corpus. NEPA experts selected nine EIS documents from different government agencies that were most representative of various NEPA actions. These documents exhibit significant variation in content and focus depending on the authoring government agency, as each agency may interpret and implement the NEPA guidelines distinctively. For instance, the U.S. Forest Service might emphasize forest management and wildfire mitigation, while the U.S. Army Corps of Engineers could prioritize water resource development and infrastructure impacts. Please refer to Table 6 in the Appendix for a brief overview about the selected documents. We consider documents of varied lengths, with the longest document containing nearly 900 pages (¿600K tokens).
@@
Gold Passage Selection For each of the nine selected documents, we attempt to select excerpts that have important content of each document. A naive approach of randomly extracting passages poses the risk of resulting in low-quality excerpts, such as parts of appendices or image captions. To avoid this risk, NEPA experts manually select excerpts from the documents. They divided each document into three sections – beginning, middle, and end, and then selected two, six, and two excerpts from each of these sections respectively, resulting a total of 10 excerpts from each document. We use these excerpts as the ground-truth context, called gold passages, for question benchmark generation.
Question Type Selection Next, we developed a question taxonomy that encompasses various complexities inherent in environmental impact assessments. After extensive discussions with NEPA experts, we finalized 10 types of questions that facilitate the NEPA decision making process. These question types test the LLM’s ability to articulate the purpose of regulations, explain complex review processes, or describe the steps involved in a permit application.
For example, funnel questions regarding the environmental alternative analysis, such as ‘Which alternatives were discussed?’, ‘Which were considered?’ ‘Why was [ALTERNATIVE] not considered?’ require in-depth environmental and legal knowledge to justify these decisions. On the other hand, problem solving questions such as ‘Given the location of the [PROJECT], create a list of aquatic species likely present in a 50-mile radius’ present specific regulatory scenarios to test the ability of the models to comprehend the latest geographical knowledge. Overall, we ensure almost half the questions in the dataset are these types of ‘open’ questions. We present the question types and relevant examples in Table 1.
Figure 2. Examples of question-answer pairs present in NEPAQuAD v1.0 , and the context used to generate these questions.
@@
In addition to selecting the question types, the NEPA experts also created sample questions for each type. For a more detailed description of the question types, as well as example questions provided by the NEPA experts, please refer to Appendices B and C.
Prompt Design We leverage generative models (such as GPT-4) with carefully crafted prompts to generate question-answer-proof triplets based on the selected EIS contexts, ensuring coverage across different question types. To ensure that the prompts can instruct generative models efficiently, we consulted with the NEPA experts to create the prompts. We also use the sample questions created by the NEPA experts to augment the original prompts and create an enhanced prompt. The template for the prompt and benchmark creation process is provided in Appendix A.
Table 1. Question types in the NEPAQuAD v1.0 benchmark
Type #Questions Type #Questions
Closed 789 49% Comparison 64 4%
Convergent 109 7% Divergent 121 8%
Evaluation 64 4% Funnel 127 8%
Inference 101 6% Problem-solving 11 1%
Process 108 7% Recall 105 7%
Automated Generation Following successful precedents in other domain specific benchmark creation (Balaguer et al., 2024), we employed GPT-4 through Azure OpenAI service with default parameters to generate the question-answer-proof benchmark triplets. Our dataset construction began with nine carefully selected EIS documents, from which we extracted 10 gold passages each, totaling 90 passages. For each passage, we generated a diverse set of questions: 10 closed-type questions and two questions each from nine other question categories. While the initial generation yielded 2520 questions, we implemented a rigorous quality control process. First, we filtered out responses that did not conform to our specified output format, reducing the count to 1920. Subsequently, we conducted manual quality assessment of the remaining questions, resulting in a final benchmark of 1590 high quality question-answer-proof triplets across the selected EIS documents. Few examples of question-answer pairs generated for a gold passage context is illustrated in Figure 2. NEPAQuAD v1.0 is made publicly available (PolicyAI, 2024) to facilitate further research and development in this domain.
2.2. Quality Assessment
Comparison with other QA datasets.  We test the intrinsic quality of the generated questions with several metrics related to content and style, and compare the metrics against the similar QA datasets SQuAD (Rajpurkar et al., 2016) and SciQ (Welbl et al., 2017). The following metrics are used to assess data quality:
  1. (1)
    Entailment between the context and the question and answer, measured via prediction from a BERT model fine-tuned on natural language inference (NLI) [^5] Accessed 1 April 2025: https://huggingface.co/tasksource/ModernBERT-base-nli;
  2. (2)
    Complexity of the question, measured via parse tree depth from a spacy dependency parser (en_core_web_trf model);
  3. (3)
    Specificity of the question, measured via number of named entities extracted using a SpaCy [^6] https://spacy.io/ named entity recognition model.
  4. (4)
    Readability of the question, measured via Flesch-Kincaid readability score (Kincaid et al., 1975).
Table 2. Comparing NEPAQuAD with existing QA benchmarks using various question quality metrics
Metric NEPAQuAD SQuAD SciQ
% Entailment + Neutral 0.925 0.985 0.968
Complexity 5.317 3.867 4.741
Specificity 1.17 1.11 0.139
Readability 34.4 60.5 58.0
The results in Table 2 show that NEPAQuAD v1.0 is on-par or exceeding in the metrics as compared to similar QA datasets. For instance, the questions in NEPAQuAD v1.0 are more complex than SQuAD and SciQ, while the proportion of questions with entailment or neutral NLI labels is in a similar range to the other datasets. The lower readability score for NEPAQuAD v1.0 indicates longer sentences and a more complex vocabulary, which is reasonable considering the NEPA domain.
Table 3. Summary of annotation agreement among three annotators.
Aspect % Majority Yes % Majority No % Unanimous Agreement
Is question type correct? 81.0 19.0 89.0
Is the answer correct? 93.0 7.0 71.0
Is the proof correct? 93.0 7.0 79.0
Table 4. Annotator agreement scores with ground truth (majority rule)
Is question type correct? Is the answer correct? Is the proof correct?
NEPA expert 1 % Agreement 99.0 81.0 87.0
Cohen’s Kappa 0.967 (Almost Perfect) 0.328 (Fair) 0.340 (Fair)
NEPA expert 2 % Agreement 90.0 90.0 92.0
Cohen’s Kappa 0.734 (Substantial) 0.423 (Moderate) 0.560 (Moderate)
Human Review To ensure a high-quality benchmark, we employed a rigorous annotation protocol. We randomly sampled 100 benchmark entries from the entire benchmark dataset. Two independent NEPA experts evaluated each sampled entry across three critical dimensions: the correctness of question type (i.e. whether the generated question was the same type of question as requested), the correctness of the generated answer, and the correctness of generated proof. For each aspect, annotators provided binary judgments (‘yes’/‘no’) based on established criteria documented in the annotation guidelines provided to the experts. For entries where the NEPA experts disagreed on one or more aspects, we engaged a third independent NEPA expert to adjudicate, creating a three-annotator subset. To determine overall quality and establish ground truth, we then employed a majority voting approach. These scores are shown in Table 3. Furthermore, we evaluated each annotator’s alignment with this majority opinion by calculating individual Cohen’s Kappa scores between each expert’s responses and the majority vote as shown in Table 4. We show examples of human annotation of the generated triplet in Appendix D.
3. The MAPLE Evaluation framework
We developed MAPLE (Multi-context Assessment Pipeline for Language model Evaluation) [^7] https://github.com/pnnl/permitai/evaluation as a comprehensive evaluation tool that streamlines the assessment of LLM s in question answering and document retrieval tasks. As illustrated in Figure 3, we designed it to provide a unified interface supporting multiple LLM providers while ensuring consistent evaluation methodologies. We incorporated a modular design that we used to integrate various context types – from no-context baselines to RAG-enhanced setups. We also implemented RAGAS evaluation metrics (Es et al., 2023), allowing us to benchmark LLM performance across different benchmarks. The MAPLE architecture comprises four modules: a flexible dataloader that standardizes benchmark entries from various file formats, a provider-agnostic LLM handler enabling unified interaction across multiple LLM services, an evaluator module that dynamically populates prompt templates with questions and context information from benchmark entries, and a comprehensive metrics module supporting both user-defined custom metrics and established metric evaluation frameworks like RAGAS (rag, 2023). We discuss the key features of MAPLE below. Please refer to Appendix G for a detailed overview of individual modules.
Figure 3. Overview of the MAPLE pipeline.
3.1. Supporting multiple LLM providers
We implemented an LLM Handler module using an abstract base class to provide support for major cloud services. Our implementation includes Azure OpenAI (GPT models), AWS Bedrock (Claude, LLama, Mistral), Google Vertex AI (Gemini), and local HuggingFace models. We designed the handler to manage provider-specific authentication requirements and parameter specifications while maintaining a standardized interface for response generation. In this way, we effectively abstract the complexities of working with different LLM services into a single, consistent evaluation pipeline.
3.2. Supporting multiple context types.
The Evaluator module (which generates LLM responses to the questions in the benchmark) supports four context types to enable comprehensive evaluation across different scenarios. These include a basic no-context mode to test prior knowledge of LLM, a PDF mode for processing complete documents, a RAG (Lewis et al., 2020) mode for vector database integration, and a gold context mode for controlled evaluations with the gold passage as context. MAPLE automatically identifies and validates appropriate context types by analyzing the available fields in benchmark entries. It defaults to the ‘no-context’ mode when only question-answer pairs are present. Here, we provide brief description of these four context types.